OpenStudy (anonymous):

Where is the vertex of the parabola? y=x^2x+3

OpenStudy (zzr0ck3r):

vertex of y=ax^2+bx+c is given by (-b/(2a), f(-b/(2a))

OpenStudy (zzr0ck3r):

so x = -b/(2a) then plug that back into f(x) to find y..

OpenStudy (zzr0ck3r):

no calc needed....

OpenStudy (anonymous):

So it's above the X-axis

OpenStudy (zzr0ck3r):

x = -b/2a = -1/2(1) = -1/2, and so y = (-1/2)^2+(-1/2) + 3 = 11/4 so (-1/2, 11/4)

OpenStudy (zzr0ck3r):

the question should say x^2+x+3?

OpenStudy (jdoe0001):

\(\bf y=x^2x+3\ ?\qquad y=x^2+3\ ?\)

OpenStudy (zzr0ck3r):

im guessing there is a missing + or -

OpenStudy (zzr0ck3r):

@Abbii_15 what should the question say?

OpenStudy (zzr0ck3r):

if it should say \(x^2+x+3\) then \(x = -b/2a = -1/2(1) = -1/2\) and so \(\\y = (-1/2)^2+(-1/2) + 3 = 11/4\) so \(\large V = (-1/2, 11/4)\)

OpenStudy (zzr0ck3r):

shes not in calc, I can tell by the other question.